# Localized Adaptive Risk Control

### v1

@@ 3.1 Electricity Demand
Figure 3: Long-term coverage (left) and average miscoverage error (right), marginalized and conditioned on weekdays and weekends for ARC and L-ARC with varying values of the localization parameter $l$ on the Elec2 dataset.
@@ 6 Acknowledgments
This work was supported by the European Union’s Horizon Europe project CENTRIC (101096379). The work of Osvaldo Simeone was also supported by the Open Fellowships of the EPSRC (EP/W024101/1) by the EPSRC project (EP/X011852/1), and by Project REASON, a UK Government funded project under the Future Open Networks Research Challenge (FONRC) sponsored by the Department of Science Innovation and Technology (DSIT).
@@ Appendix A Proof of Theorem 1
As stated in Lemma 1, we can always find $T$ large enough, such that $\left\lVert g^{*}-\bar{g}_{T}\right\rVert_{\infty}\leq\epsilon$ with arbitrary large probability. This implies, that for any $\delta^{\prime}>0$,
$\lim_{T\to\infty}\max_{w\in\mathcal{W}}\mathbb{E}\left[\frac{w(X)}{\mathbb{E}[w(X)]}\mathcal{L}\left(C\left(X,\bar{g}_{T}\right),Y\right)\right]$ $\leq\max_{w\in\mathcal{W}}\mathbb{E}\left[\frac{w(X)}{\mathbb{E}[w(X)]}\mathcal{L}\left(C\left(X,g^{*}\right),Y\right)\right]+\delta^{\prime}$ (55)
@@ C.2 Image Classification with Calibration Requirements
In the leftmost panel of Figure 8 we report the long-term coverage of ARC and L-ARC for an increasing level of localization obtained by decreasing the length scale $l$. All methods guarantee long-term coverage. In the middle panel, we use hold-out data to evaluate the coverage of the calibrated model conditioned on the model’s confidence level. For small length scale $l$, L-ARC yields prediction sets that satisfy the coverage requirement across different levels of the model’s confidence. In contrast, ARC, due to its inability to adapt the threshold function, has a large miscoverage rate for small model confidence levels. As illustrated in the right panel, this is achieved by producing a larger set size when the model’s confidence is low.

### v3

@@ 2.2 L-ARC
It is important to note that, at any time $t$, computing the threshold function (19) requires storing the coefficients $\{a^{i}_{t}\}^{t-1}_{i=1}$ and $c_{t}$, as well as the input data $\{X_{t}\}^{t}_{i=1}$. Consequently, L-ARC has a linear memory requirement in $t$, which is a known limitation of non-parametric learning in online settings (Koppel et al., 2020). Previous research has explored methods that trade memory efficiency for accuracy (Kivinen et al., 2004). In Appendix C.3, we build on these approaches to present a memory-efficient variant of L-ARC that allows for a trade-off between localized risk control and memory requirements.
@@ 3.1 Electricity Demand
Figure 3: Long-term coverage (left) and average miscoverage error (right), marginalized and conditioned on weekdays and weekends. for ARC and L-ARC with varying values of the localization parameter $l$ on the Elec2 dataset.
@@ 6 Acknowledgments
This work was supported by the European Union’s Horizon Europe project CENTRIC (101096379). The work of Osvaldo Simeone was also supported by the Open Fellowships of the EPSRC (EP/W024101/1) by the EPSRC project (EP/X011852/1), and by Project REASON, a UK Government funded project under the Future Open Networks Research Challenge (FONRC) sponsored by the Department of Science Innovation and Technology (DSIT). We would also like to express our gratitude to Anastasios Angelopoulos for valuable insights on the technical content of the paper.
@@ Appendix A Proof of Theorem 1
As stated in Lemma 1, we can always find $T$ large enough, such that $\left\lVert g^{*}-\bar{g}_{T}\right\rVert_{\infty}\leq\epsilon$ with arbitrary large probability. This implies, that for any $\delta^{\prime}>0$ and $w\in\mathcal{W}$,
$\lim_{T\to\infty}\mathbb{E}\left[\frac{w(X)}{\mathbb{E}[w(X)]}\mathcal{L}\left(C\left(X,\bar{g}_{T}\right),Y\right)\right]$ $\leq\mathbb{E}\left[\frac{w(X)}{\mathbb{E}[w(X)]}\mathcal{L}\left(C\left(X,g^{*}\right),Y\right)\right]+\delta^{\prime}$ (55)
@@ C.2 Image Classification with Calibration Requirements
In the left-most panel of Figure 8 we report the long-term coverage of ARC and L-ARC for an increasing level of localization obtained by decreasing the length scale $l$. All methods guarantee long-term coverage. In the middle panel, we use hold-out data to evaluate the coverage of the calibrated model conditioned on the model’s confidence level. For small length scale $l$, L-ARC yields prediction sets that satisfy the coverage requirement across different levels of the model’s confidence. In contrast, ARC, due to its inability to adapt the threshold function, has a large miscoverage rate for small model confidence levels. As illustrated in the right panel, this is achieved by producing a larger set size when the model’s confidence is low.
C.3 On the memory efficiency of L-ARC
In a manner similar to [Kivinen et al., 2004], it is possible to obtain a memory-efficient version of L-ARC that adopts a truncated version of L-ARC threshold (19) given by
$g_{t+1}(\cdot)=\sum^{t}_{i=\max\{1,t-M_{\text{max}}\}}a^{i}_{t+1}k(X_{i},\cdot)+c_{t+1}.$ (86)
Unlike the threshold (19), which has a linear memory requirement, the truncated version (86) requires a constant memory and computational load that are proportional to the number of coefficients $M_{\text{max}}$. It is known that in online non-parametric learning, there exists a trade-off between memory efficiency and performance. In the following, we empirically study the trade-off between the localized risk control of L-ARC and its memory requirements by varying the parameteR $M_{\text{max}}$.
C.3.1 Tumor Segmentation
Figure 9: FNR obtained by ARC, L-ARC, and L-ARC with limited memory budget $M_{\text{max}}\in\{500,1000,1500\}$. As the memory budget increases, the localized risk control performance of L-ARC interpolates between ARC and L-ARC.
Using the setup described in Section 3.2, we now consider calibrating the image segmentation model using L-ARC with a truncated threshold (86). In Figure 9, we report the average FNR conditioned on different data sources for $M_{\text{max}}\in\{500,1000,1500\}$. As a benchmark, we also compare against ARC and L-ARC without truncation. By adjusting the value of $M_{\text{max}}$, it is possible to trade off localized risk control for memory efficiency. In fact, the effect of truncation on L-ARC’s performance is minimal when the number of coefficients in the truncation is large $(M_{\text{max}}=1500)$. However, for greater memory savings $(M_{\text{max}}=500)$, L-ARC’s performance becomes similar to that of ARC. In all cases, L-ARC provides better localized risk control than ARC.
C.3.2 Beam Selection
Figure 10: SNR across the deployment attained by L-ARC with limited memory budget $M_{\text{max}}$.
We consider the beam selection problem discussed in Section 3.3. In Figure 10, we report the SNR levels across the deployment attained by ARC, L-ARC, and L-ARC with a truncated threshold with a maximum number of coefficients $M_{\text{max}}\in\{500,1000\}$. As the number of coefficients $M_{\text{max}}$ and the memory requirement reduce, the localized risk control performance of L-ARC also decreases. Nonetheless, even for small $M_{\text{max}}$, L-ARC delivers a more consistent SNR level across the deployment compared to ARC.